# Global Optimal Closed-Form Solutions for Intelligent Surfaces With Mutual Coupling: Is Mutual Coupling Detrimental or Beneficial?

### v1

@@ VI Impact of Mutual Coupling
Given a positive definite matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ with diagonal elements $[\mathbf{A}]_{n,n}=a$, for $n=1,\ldots,N$, it holds
@@
Given a positive definite matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ with diagonal elements $[\mathbf{A}]_{n,n}=a$, for $n=1,\ldots,N$, it holds
@@
By using the results in Lemmas 88 and 89, we can derive the following proposition, stating that mutual coupling is beneficial as it improves the average channel gain.
@@ VII Numerical Results
where $\eta_{0}=377\;\Omega$ is the impedance of free space, $k_{0}=2\pi/\lambda$ is the wavenumber, and $d_{q,p}=\sqrt{(x_{q}-x_{p})^{2}+(y^{\prime\prime}-y^{\prime})^{2}}$. We consider the direct link between the transmitter and receiver to be fully obstructed, i.e., $z_{RT}=0$, and generate $\mathbf{z}_{RI}$ and $\mathbf{z}_{IT}$ as independent Rayleigh distributed, i.e., $\mathbf{z}_{RI}\sim\mathcal{CN}(\mathbf{0},\rho_{RI}\mathbf{I})$ and $\mathbf{z}_{IT}\sim\mathcal{CN}(\mathbf{0},\rho_{IT}\mathbf{I})$, with path gains $\rho_{RI}=\rho_{IT}=4Z_{0}^{2}10^{-8}$.
@@
Figure 5: Channel gain versus the inter-element distance $d$ achieved by BD-RIS and D-RIS reconfigured through mutual coupling-aware and -unaware optimization. In the case of coupling-unaware optimization, we assume the channels in the $Z$-parameters $\mathbf{z}_{RI}$, $\mathbf{z}_{IT}$, and $z_{RT}$ to be known.
@@ VII-B Scaling Laws
In Section V, we have studied the mean value of the channel gain upper bounds (51) and (75), i.e., their scaling laws, which are provided in closed-form in (74) and (87), respectively. We now verify the accuracy of these scaling laws. To this end, in Fig. 3, we individually analyze the expectation terms appearing in (52) and (76), reporting their simulated mean and the theoretical mean derived in closed-form in Section V. We observe that the theoretical mean values are accurate for practical numbers of RIS elements, such as $N_{I}=64$, even if derived with approximations valid at $N_{I}\rightarrow\infty$.
@@ VII-C Benefits of BD-RIS and Mutual Coupling-Aware Optimization
  • For D-RIS optimized through a mutual-coupling aware method, we consider the optimization method proposed in [15].
@@
We make the following observations from Fig. 5. First, when optimizing a BD-RIS or D-RIS accounting for mutual coupling, the performance increases with the mutual coupling strength, i.e., as the inter-element distance decreases. However, when BD-RIS or D-RIS are optimized not being aware of the mutual coupling, their performance dramatically drops for short inter-element distances, experiencing a degradation of up to 5 dB for BD-RIS and 4 dB for D-RIS. This is because the RIS is optimized based on a channel model with no mutual coupling, which differs from the physics-consistent channel model including the mutual coupling effects. Second, the gain of BD-RIS over D-RIS is approximately 2 dB when optimized accounting for mutual coupling, which is $1.6$ times in linear scale, consistently with the gain derived in [3] in the absence of mutual coupling. This gain slightly increases with the mutual coupling strength, i.e., as the inter-element distance decreases, since BD-RIS offers additional flexibility to deal with mutual coupling. Third, the observed trends of the channel gain versus the inter-element distance are independent of the number of RIS elements, as it can be seen by comparing Fig. 5(a) and 5(b), where we fix $N_{I}=64$ and $N_{I}=128$, respectively.
@@ VIII-A Proof of Proposition 1
where we exploited the fact that $\mathbf{Z}$ is symmetric for reciprocal networks. Since $\mathbf{Z}$ is symmetric, we have $[\mathbf{Z}]_{m,n}=[\mathbf{Z}]_{n,m}$, yielding
@@ VIII-B Proof of Lemma 88
which is $\text{Tr}\left(\mathbf{A}\right)=aN$ since $[\mathbf{A}]_{n,n}=a$, for $n=1,\ldots,N$. Besides, $\lambda_{1}^{-1},\ldots,\lambda_{N}^{-1}$ are the $N$ eigenvalues of $\mathbf{A}^{-1}$, whose sum is given by
@@ VIII-D Proof of Proposition 2
which can be done by applying Lemmas 88 and 89. The inequality in (108) can be proved by applying Lemma 89 with $\mathbf{A}=\Re\{\mathbf{Z}_{II}\}$ and $a=Z_{II}$. Note that $\Re\{\mathbf{Z}_{II}\}$ fulfills the hypothesis of Lemma 89 since it is positive definite following Proposition 1. Besides, the inequality in (109) can be proved by proving $\text{Tr}(\Re\{\mathbf{Z}_{II}\}^{-1})\geq N/Z_{II}$, which follows from Lemma 88. Finally, the inequalities in (110) and (111) directly follows from (108) and (109), respectively, concluding the proof that $\mathbb{E}[|h_{MC}^{\star}|^{2}]\geq\mathbb{E}[|h_{NoMC}^{\star}|^{2}]$.

### v2

@@ VI Impact of Mutual Coupling
Given a positive definite matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ with diagonal elements $\mathbf{A}]_{n,n}=a$, for $n=1,\ldots,N$, it holds
@@
Given a positive definite matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ with diagonal elements $\mathbf{A}]_{n,n}=a$, for $n=1,\ldots,N$, it holds
@@
By using the results in Lemmas 91 and 92, we can derive the following proposition, stating that mutual coupling is beneficial as it improves the average channel gain.
@@ VII Numerical Results
where $\eta_{0}=377\;\Omega$ is the impedance of free space, $k_{0}=2\pi/\lambda$ is the wavenumber, and $d_{q,p}=\sqrt{(x_{q}-x_{p})^{2}+(y^{\prime\prime}-y^{\prime})^{2}}$. We consider the direct link between the transmitter and receiver to be fully obstructed, i.e., $z_{RT}=0$, and generate $\mathbf{z}_{RI}$ and $\mathbf{z}_{IT}$ as independent Rayleigh distributed, i.e., $\mathbf{z}_{RI}\sim\mathcal{CN}(\mathbf{0},\rho_{RI}\mathbf{I})$ and $\mathbf{z}_{IT}\sim\mathcal{CN}(\mathbf{0},\rho_{IT}\mathbf{I})$, with path gains $\rho_{RI}=\rho_{IT}=4Z_{0}^{2}10^{-8}$. Although the path gains $\rho_{RI}$ and $\rho_{IT}$ depend on the transmission distances, we directly set the value of $\rho_{RI}$ and $\rho_{IT}$ since they impact the channel gain by just scaling it by the product $\rho_{RI}\rho_{IT}$.
@@
Figure 6: Channel gain versus the inter-element distance $d$ achieved by BD-RIS and D-RIS reconfigured through mutual coupling-aware (denoted as “MC aware”) and -unaware (denoted as “MC unaware”) optimization. In the case of coupling-unaware optimization, we assume the channels in the $Z$-parameters $\mathbf{z}_{RI}$, $\mathbf{z}_{IT}$, and $z_{RT}$ to be known.
@@ VII-B Scaling Laws
In Section V, we have studied the mean value of the channel gain upper bounds (54) and (78), i.e., their scaling laws, which are provided in closed-form in (77) and (90), respectively. We now verify the accuracy of these scaling laws. To this end, in Fig. 4, we individually analyze the expectation terms appearing in (55) and (79), reporting their simulated mean and the theoretical mean derived in closed-form in Section V. We observe that the theoretical mean values are accurate for practical numbers of RIS elements, such as $N_{I}=64$, even if derived with approximations valid at $N_{I}\rightarrow\infty$. Additional numerical results have shown that the scaling laws remain accurate for a number of RIS elements as low as $N_{I}=16$.
@@ VII-C Benefits of BD-RIS and Mutual Coupling-Aware Optimization
  • For D-RIS optimized through a mutual-coupling aware method, we consider the optimization method proposed in [15], which has a computational complexity of $\mathcal{O}(IN_{I}^{3})$, where $I$ is the number of iterations needed to converge.
@@
We make the following observations from Fig. 6. First, when optimizing a BD-RIS or D-RIS accounting for mutual coupling, the performance increases with the mutual coupling strength, i.e., as the inter-element distance decreases. However, when BD-RIS or D-RIS are optimized not being aware of the mutual coupling, their performance dramatically drops for short inter-element distances, experiencing a degradation of up to 5 dB for BD-RIS and 4 dB for D-RIS. This is because the RIS is optimized based on a channel model with no mutual coupling, which differs from the physics-consistent channel model including the mutual coupling effects. Second, the gain of BD-RIS over D-RIS is approximately 2 dB when optimized accounting for mutual coupling, which is $1.6$ times in linear scale, consistently with the gain derived in [3] in the absence of mutual coupling. This gain slightly increases with the mutual coupling strength, i.e., as the inter-element distance decreases, since BD-RIS offers additional flexibility to deal with mutual coupling. Third, the observed trends of the channel gain versus the inter-element distance are independent of the number of RIS elements, as it can be seen by comparing Fig. 6(a) and 6(b), where we fix $N_{I}=64$ and $N_{I}=128$, respectively. Fourth, the impact of mutual coupling can be approximately neglected when the inter-element distance is larger than half-wavelength, i.e., $\lambda/2$, and the RIS can be successfully optimized without considering mutual coupling in this case.
@@ VIII Conclusion
Future research directions include the extension of our results to multi-antenna and multi-user systems, which is challenging since closed-form solutions and scaling laws are often not known even in the absence of mutual coupling. In addition, more realistic studies could analyze RIS in the presence of mutual coupling jointly with other hardware impairments and non-idealities, such as losses, discrete-value tunable components, and imperfect CSI. Experimental validation is needed to verify whether the derived performance upper bound can be achieved in real-world scenarios.
@@ VIII-A Proof of Proposition 1
where we exploited the fact that $\mathbf{Z}$ is symmetric for reciprocal networks. Since $\mathbf{Z}$ is symmetric, we have $\mathbf{Z}]_{m,n}=[\mathbf{Z}]_{n,m}$, yielding
@@ VIII-B Proof of Lemma 91
which is $\text{Tr}\left(\mathbf{A}\right)=aN$ since $\mathbf{A}]_{n,n}=a$, for $n=1,\ldots,N$. Besides, $\lambda_{1}^{-1},\ldots,\lambda_{N}^{-1}$ are the $N$ eigenvalues of $\mathbf{A}^{-1}$, whose sum is given by
@@ VIII-D Proof of Proposition 2
which can be done by applying Lemmas 91 and 92. The inequality in (111) can be proved by applying Lemma 92 with $\mathbf{A}=\Re\{\mathbf{Z}_{II}\}$ and $a=Z_{II}$. Note that $\Re\{\mathbf{Z}_{II}\}$ fulfills the hypothesis of Lemma 92 since it is positive definite following Proposition 1. Besides, the inequality in (112) can be proved by proving $\text{Tr}(\Re\{\mathbf{Z}_{II}\}^{-1})\geq N/Z_{II}$, which follows from Lemma 91. Finally, the inequalities in (113) and (114) directly follows from (111) and (112), respectively, concluding the proof that $\mathbb{E}[|h_{MC}^{\star}|^{2}]\geq\mathbb{E}[|h_{NoMC}^{\star}|^{2}]$.